OpenStudy (anonymous):

108 is 36% of what number? Write and solve a proportion to solve the problem

OpenStudy (anonymous):

108 is 36 of what number. 108 divided by 0.36 = 300

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Welcome

OpenStudy (mathstudent55):

100% corresponds to the number 36% corresponds to a part which is 108 100% is to x as 36% is to 108 100/x = 36/108 That's the proportion. Now cross multiply and solve for x

OpenStudy (anonymous):

360?

OpenStudy (anonymous):

Wait what.

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

100/x = 36/108 Cross multiply: 36x = 100(108) 36x = 18,800 x = 300 Which is what Lilmoney got above.

OpenStudy (anonymous):

thanks :

OpenStudy (mathstudent55):

You're welcome